# Experimental Design for Active Transductive Inference in Large Language Models

### v1

@@ 3.2 Simulation-Based Algorithm
1:  Input: Target point $\mathbf{x}_{*}$, budget $T$
2:
3:  $\mathcal{L}_{1}\leftarrow\emptyset,\ \mathcal{U}_{1}\leftarrow[n],\ H_{1}\leftarrow()$
4:  for $t=1,\dots,T$ do
5:     for all $i\in\mathcal{U}_{t}$ do
6:        for all $j\in[m]$ do
7:           Sample $Y_{t,i,j}\sim p(\cdot\mid\mathbf{x}_{i},H_{t})$
8:           $H_{t,i,j}\leftarrow H_{t}\oplus(\mathbf{x}_{i},Y_{t,i,j})$
9:           Sample $\tilde{Y}_{t,i,j,1},\tilde{Y}_{t,i,j,2}\sim p(\cdot\mid\mathbf{x}_{*},H_{t,i,j})$
@@
12:     $I_{t}\leftarrow\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})^{2}$
13:     $X_{t}\leftarrow\mathbf{x}_{I_{t}}$
14:     Observe label $Y_{t}$ of point $X_{t}$
15:     $\mathcal{L}_{t+1}\leftarrow\mathcal{L}_{t}\cup\{I_{t}\},\ \mathcal{U}_{t+1}\leftarrow\mathcal{U}_{t}\setminus\{I_{t}\}$
16:     $H_{t+1}\leftarrow H_{t}\oplus(X_{t},Y_{t})$
17:  end for
18:
19:  Output: Sample $Y_{*}\sim p(\cdot\mid\mathbf{x}_{*},H_{T+1})$
@@
The shortcoming of GO is that the chosen point in round $t$, $X_{t}$, is independent of the previously observed labels, because (1) does not depend on $(Y_{\ell})_{\ell\in[t-1]}$. This is a property of linear models that is undesirable in non-linear models. This motivates our second algorithm design, which simulates the impact of labeling points on uncertainty. Therefore, we call it Simulation-Based Active Learning and abbreviated it as SAL. The pseudocode of SAL is in Algorithm 2.
The key idea in our approach is to replace the closed-form formula in (1) by a simulation. Fix round $t$, history $H_{t}$, and a candidate point $\mathbf{x}_{i}$. To estimate its impact, we simulate its labels $m$ times, by sampling them as $Y_{t,i,j}\sim p(\cdot\mid\mathbf{x}_{i},H_{t})$ for all $j\in[m]$. Let
$H_{t,i,j}=H_{t}\oplus(\mathbf{x}_{i},Y_{t,i,j})$
be the history $H_{t}$ extended by point $\mathbf{x}_{i}$ and its simulated label $Y_{t,i,j}$. Then, to estimate the reduction in uncertainty of $Y_{*}$, we take two independent samples
$\tilde{Y}_{t,i,j,1},\tilde{Y}_{t,i,j,2}\sim p(\cdot\mid\mathbf{x}_{*},H_{t,i,j})$
for each $j\in[m]$ and score point $\mathbf{x}_{i}$ by
$\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})^{2}\,.$ (2)
The point with the lowest score is chosen and we denote its index by $I_{t}$.
@@
This approach can be justified as follows. Take the same setting as in Section 3.1. Given label $Y_{t,i,j}$ for point $\mathbf{x}_{i}$, the posterior distribution of $\bm{\theta}_{*}\mid H_{t,i,j}$ is $\mathcal{N}(\widehat{\bm{\theta}}_{t,i,j},\widehat{\mathbf{\Sigma}}_{t,i})$, where
$\widehat{\mathbf{\Sigma}}_{t,i}=(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\sigma^{-2}\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}$
is the simulated posterior covariance of $\bm{\theta}_{*}$ and
$\widehat{\bm{\theta}}_{t,i,j}=\widehat{\mathbf{\Sigma}}_{t,i}\left(\widehat{\mathbf{\Sigma}}_{0}^{-1}\bm{\theta}_{0}+\sigma^{-2}\left(\sum_{\ell=1}^{t-1}X_{\ell}Y_{\ell}+\mathbf{x}_{i}Y_{t,i,j}\right)\right)$
is the corresponding posterior mean. By design, $\tilde{Y}_{t,i,j,1}$ and $\tilde{Y}_{t,i,j,2}$ are independent samples from $\mathcal{N}(\mathbf{x}_{*}^{\top}\bm{\theta}_{*},\sigma^{2})$, where $\bm{\theta}_{*}\sim\mathcal{N}(\widehat{\bm{\theta}}_{t,i,j},\widehat{\mathbf{\Sigma}}_{t,i})$. Therefore,
$\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2}\sim\mathcal{N}(0,2(\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}))\,.$
This distribution does not depend on $Y_{t,i,j}$. It also follows that (2) is an unbiased estimate of $2(\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2})$, proportional to $\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}$ in (1). Thus, in linear models, SAL can be viewed as an inefficient implementation of GO.
@@
The time complexity of SAL is $O(nmT)$. This is because the algorithm searches for the best point out of at most $n$ points in $T$ rounds. The evaluation of each point in each round requires $3m$ LLM queries. Therefore, although the time complexity of SAL seems much smaller than that of GO, SAL is much slower in practice.
3.3 Vector Labels
@@
GO and SAL can be easily extended to vector labels. GO does not depend on labels at all. The only modification in SAL is that (2) is replaced with
$\frac{1}{m}\sum_{j=1}^{m}\|\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})\|_{2}^{2}\,.$ (3)
This is the sum of the posterior variances of label $Y_{*}$ for point $\mathbf{x}_{i}$, over all dimensions of $Y_{*}$.
3.4 Multiple Target Points
We extend GO and SAL to $K$ target points. As a result, they can be applied to classification and regression tasks, where we may want to infer labels of multiple points. We adopt the following notation. The $k$-th target point is $\mathbf{x}_{*,k}$ and its label is $Y_{*,k}$. GO can be extended as follows. In round $t$, the chosen point minimizes the maximum uncertainty of predictions over all target points, which corresponds to
$I_{t}=\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*,k}\,.$
SAL is modified as follows. Let $\tilde{Y}_{t,i,j,k,1}$ and $\tilde{Y}_{t,i,j,k,2}$ be simulated labels of the LLM for target point $\mathbf{x}_{*,k}$ in round $t$. Then we replace (3) with
$\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\|\tilde{Y}_{t,i,j,k,1}-\tilde{Y}_{t,i,j,k,2})\|_{2}^{2}\,.$
Finally, both GO and SAL predict labels $Y_{*,k}$ for all target points $\mathbf{x}_{*,k}$ at the end of round $T$.

### v2

@@ 3.2 Simulation-Based Algorithm
1:  Input: Training set $\mathcal{X}_{\text{examples}}=\{\mathbf{x}_{i}\}_{i=1}^{n}$, test set $\mathcal{X}_{*}=\{\mathbf{x}_{*,k}\}_{k=1}^{K}$, budget $T$
2:  $\mathcal{L}_{1}\leftarrow\emptyset,\ \mathcal{U}_{1}\leftarrow[n],\ H_{1}\leftarrow\{\}$
3:  for $t=1,\dots,T$ do
4:     for all $i\in\mathcal{U}_{t}$ do
5:        for all $\mathbf{x}_{*,k}\in\mathcal{X}_{*}$ do
6:           for $j=1,2,\ldots,m$ do
7:              Sample $Y^{(j)}_{t,i}\sim p(\cdot\mid\mathbf{x}_{i},H_{t})$
8:              $H_{t,i,j}\leftarrow H_{t}\cup\{(\mathbf{x}_{i},Y^{(j)}_{t,i})\}$
9:              Sample $\tilde{Y}^{(j,1)}_{t,i,k},\tilde{Y}^{(j,2)}_{t,i,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{t,i,j})$
@@
12:     end for
13:     $I_{t}\leftarrow\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\left(\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k}\right)^{2}$
14:     $X_{t}\leftarrow\mathbf{x}_{I_{t}}\in\mathcal{X}_{\text{examples}}$
15:     Observe label $Y_{t}$ of example $X_{t}$
16:     $\mathcal{L}_{t+1}\leftarrow\mathcal{L}_{t}\cup\{I_{t}\},\ \mathcal{U}_{t+1}\leftarrow\mathcal{U}_{t}\setminus\{I_{t}\}$
17:     $H_{t+1}\leftarrow H_{t}\cup\{(X_{t},Y_{t})\}$
18:  end for
19:  Output: Sample $Y_{*,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{T+1})$ for $k\in[K]$
@@
While GO reduces uncertainty in label predictions, it has a major limitation. The chosen example $X_{t}$ at round $t$ is not affected by observed labels $(Y_{\ell})_{\ell\in[t-1]}$. This is because (1) does not depend on $(Y_{\ell})_{\ell\in[t-1]}$. While this is a property of linear models, it is undesirable in non-linear models, such as LLMs. To address this limitation, we propose a new algorithm that simulates the impact of labeling examples on the uncertainty of predicted labels. We call it Simulation-Based Active Learning and abbreviated it as SAL. The pseudocode of SAL is provided in Algorithm 2.
The key idea in SAL is to replace the closed-form formula in (1) by a simulation. We detail the algorithm next. Fix round $t$, history $H_{t}$, and candidate example $\mathbf{x}_{i}$. To estimate the impact of labeling $\mathbf{x}_{i}$, we simulate its labels $m$ times. For each simulation $j\in[m]$, we sample $Y^{(j)}_{t,i}$ from the conditional distribution $p(\cdot\mid\mathbf{x}_{i},H_{t})$ using the LLM. Then we extend the history $H_{t}$ by $\mathbf{x}_{i}$ and its simulated label $Y^{(j)}_{t,i}$, $H_{t,i,j}=H_{t}\cup\{(\mathbf{x}_{i},Y^{(j)}_{t,i})\}$. This process results in $m$ copies of augmented histories, each reflecting a potential outcome of labeling of $\mathbf{x}_{i}$. Finally, we take two independent samples for each $j\in[m]$ as $\tilde{Y}^{(j,1)}_{t,i,k},\tilde{Y}^{(j,2)}_{t,i,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{t,i,j})$. The maximum uncertainty over test examples after labeling $\mathbf{x}_{i}$ is estimated as
$\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\left(\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k}\right)^{2}\,.$ (2)
The training example with the lowest value is chosen and we denote its index by $I_{t}$. Then $\mathbf{x}_{I_{t}}$ and its observed label $Y_{t}$ are added to the history to get $H_{t+1}$ for the next iteration $t+1$.
@@
Next we justify SAL. Consider the same setting as in Section 3.1. Given a label $Y^{(j)}_{t,i}$ for example $\mathbf{x}_{i}$, the posterior distribution of $\bm{\theta}_{*}\mid H_{t,i,j}$ is $\mathcal{N}(\widehat{\bm{\theta}}_{t,i,j},\widehat{\mathbf{\Sigma}}_{t,i})$, where $\widehat{\mathbf{\Sigma}}_{t,i}=(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\sigma^{-2}\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}$ is the simulated posterior covariance of $\bm{\theta}_{*}$ and
$\widehat{\bm{\theta}}_{t,i,j}=\widehat{\mathbf{\Sigma}}_{t,i}\left(\widehat{\mathbf{\Sigma}}_{0}^{-1}\bm{\theta}_{0}+\sigma^{-2}\left(\sum_{\ell=1}^{t-1}X_{\ell}Y_{\ell}+\mathbf{x}_{i}Y^{(j)}_{t,i}\right)\right)$
[…]
@@
The time complexity of SAL is $O(nKmT)$. This is because it searches for the best example out of at most $n$ in $T$ rounds for each test example $k\in[K]$. The evaluation of impact on each test example requires $2m$ LLM queries.
@@
Vector labels: GO and SAL are easy to extend to vector labels, $d_{y}>1$. GO does not depend on labels at all. The only modification in SAL is that (2) is replaced with $\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\|\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k})\|_{2}^{2}$. This is the sum of the posterior variances of the labels over all dimensions.